OpenStudy (anonymous):

simplify 4.2(x+y)+3.3(x+y)+5x

OpenStudy (callisto):

4.2(x+y)+3.3(x+y)+5x =4.2x + 4.2 y + 3.3x + 3.3y + 5x = 4.2x + 3.3x + 5x + 4.2y + 3.3y = (4.2 + 3.3 + 5)x + (4.2 + 3.3) y = ... ?

OpenStudy (callisto):

By the way, is it \(4.2\) or \(4 \cdot 2\)?